OpenStudy (anonymous):

FIRST CORRECT ANSWER GETS A MEDAL! Reese read twice as many pages Saturday night than she read Sunday night. If she read a total of 78 pages over the weekend, how many pages did Reese read Sunday night?

OpenStudy (gorv):

we have no idea how much pages she read on sunday night

OpenStudy (gorv):

so let she read x pages on sunday

OpenStudy (gorv):

on saturday she read twice of sunday so page read= 2*x

OpenStudy (gorv):

you got upto here??

OpenStudy (anonymous):

yea I think so

OpenStudy (gorv):

if u have any doubt ask ???

OpenStudy (gorv):

we know that she read total 78 pages on sat n sun

OpenStudy (gorv):

total=78=saturday pages +sunday pages 78=x+2x 78=3x

OpenStudy (gorv):

x=????

OpenStudy (gorv):

ypou solve and tell me ????

OpenStudy (anonymous):

26?

OpenStudy (gorv):

yuppppppppppppppppppppppp

OpenStudy (anonymous):

thank you

OpenStudy (gorv):

welcome buddy :)